OpenStudy (anonymous):

A line contains the points (-1,k), (6,k-1), and (20,3). What is the value of k?

OpenStudy (anonymous):

u can answer, but i have to go eat dinner right now so please don't feel like I'm not listening. Please answer if you know, will give medal

jimthompson5910 (jim_thompson5910):

Label the 3 points A,B,C So A = (-1,k) B = (6,k-1) C = (20,3) --------------------------------------------------------- Step 1) find the expression for the slope of the line through A and C \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\Large m = \frac{3-k}{20-(-1)}\] \[\Large m = \frac{3-k}{20+1}\] \[\Large m = \frac{3-k}{21}\] So the slope expression for the line through A and C is \(\Large \frac{3-k}{21}\) --------------------------------------------------------- Step 2) find the expression for the slope of the line through B and C \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\Large m = \frac{3-(k-1)}{20-6}\] \[\Large m = \frac{3-k+1}{14}\] So the slope expression for the line through B and C is \(\Large \frac{3-k+1}{14}\) --------------------------------------------------------- Step 3) equate the two slope expressions from step 1) and step 2) \[\Large \frac{3-k}{21}=\frac{3-k+1}{14}\] I'll let you solve for k

OpenStudy (anonymous):

okay thank you very much @jim_thompson5910

OpenStudy (anonymous):

you always give the best explanations and answers

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@jim_thompson5910 does k = 6?

jimthompson5910 (jim_thompson5910):

yep you nailed it

OpenStudy (anonymous):

not really, it took me too long, but thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome